OpenStudy (anonymous):

f(x)=x^2-x-1, g(x)=square root x-5 how do i find (f o g)

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(\sqrt{x-5})\] is a start

OpenStudy (anonymous):

then since \[f(\diamondsuit)=\diamondsuit^2-\diamondsuit-1\] you have \[f(\sqrt{x-5})=(\sqrt{x-5})^2-\sqrt{x-5}-1\]